Two of the eigenvalues of a 3*3 matrix A are 2 and 3 and the determinant is 48. Find the third eigenvalue and the trace (A)


1
Expert's answer
2021-07-12T15:01:40-0400

Determinant is the product of eigenvalues


"\det A=\lambda_1\lambda_2\lambda_3=48"

Given "\lambda_1=2, \lambda_2=3." Then


"\lambda_3=\dfrac{\det A}{\lambda_1\lambda_2}=\dfrac{48}{2(3)}=8"

"tr(A)=\lambda_1+\lambda_2+\lambda_3=2+3+8=13"